Trouble Shooting


Table of Contents

Can't find fringes

Restarting Servers

Telescopes and Dome Servers

Tiptilt Server

OPLE and Metrology

Beam Samplers

Remote Observing

General Problems

This document describes how to solve common problems that sometimes arise at the array.


Can't find fringes

This is a list of simple things to check if you can't find fringes:
  • Were the clocks synced? Make sure the [SYNC CLOCKS] button on Cosmic Debris has been pushed to start the night. If the OPLE server does not display the correct CHARA time and the errors don't read (0) or (1), the clocks were not synced. As the VME runs on its own clock and does not use the NTP server, it can drift over the course of the night and cause problems with finding fringes, even if it was synced at the start of the night. Syncing the clocks multiple time during the night may help to avoid this hidden clock problem.
  • Did the Astrolib update on OPLE? If the job queue is stopped too soon after slewing on Cosmic Debris, the correct calculations for the carts will not be done by OPLE and you may be searching for fringes with the wrong star data. The proper star can be entered by typing hd .... into the OPLE server and hitting ENTER.
  • Is the target a high proper motion star? Red dwarfs are close stars and can have high proper motions. Scan a wider range to see if it is outside of the usual calculated scan range. Binaries can also have very high offsets from the expected position due to mistakenly using astromod calculations from the companion star.
  • Did you get the same star in each telescope? Sometimes a busy star field and poor pointing of the telescopes can lead to the wrong star being acquired and locked by tiptilt. View the stars in the finder window to see if all the stars match.
  • Check the CHARA time on the GPS server. The "Ext-CHARA," "CHARA-Sys," and "Ext-Sys" time offsets listed on the GPS server should be small (< 0.01 sec). If there are large time offsets, then a GSYNC might be needed.
  • Check the time on the OPLE server. If the time is off or there are any lost ticks/seconds "Lost T/S" listed, then MSYNC ople using the GPS GUI and type "syncople" into the ople server.
  • Are the [MAN] buttons pressed (gray) for the moving carts on the OPLE Control gui? (The reference cart will remain green.)
  • Check that the carts are within delay line range (-1 to 44 meters) and errors are small. Were the carts homed before the first slew of the night? Did any carts go to the front switch after slewing? This can cause them to lose their position. Turn off the OL button to send it back to zero. If it does not track on the home switch when displaying a zero position (0.000000), it was lost. Is the E2 cart stuck at 27.97m? How are the metrology signals? Are they strong and staying white? Red signals mean one or more metrology signals have gone too low and homing carts is necessary. See section on homing carts in OPLE and Metrology section below.
  • Are the LDC's working?
    • Is the glass position within allowable range (-10mm to 49mm) on all beams?
    • Have the commands "useldc on" and "autoldc on" been typed into the ople server?
    • Look out for "-10" and "49" errors on ople server. If the LDC glass goes out of range, the LDC will need to be homed. To home the glass, stop astromod by clicking [STOP] on the OPLE gui. Make sure the LDC velocities are non-zero (~ 50... must click [VEL] button to update velocity) and click [HOME] on the LDCs in use. The LDC position number should go to 0.000 or very close to zero. After the LDCs have been homed, turn astromod on by clicking [START] on the ople gui. Then click [REF] on cosmic debris to send the LDCs to the correct positions.
    • If LDCs are working correctly, the LDC positions (in mm) should roughly match up with the reference cart positions (in meters).

  • Check the instrument alignment. Is flux getting through to the detector? How long has it been since the last NIRO camera alignment? Classic and CLIMB programs can run for about an hour before the light will drift from the central pixel. Use the Classic or CLIMB gui to view the light on the pixels by clicking the PICTURE tab and then the PIXEL AREA button. Turn the camera off after a few seconds to avoid crashing NIRO. Is the right dither power turned on? CLIMB 1 and Classic use different dithers. If Classic or CLIMB fringes are found in a scan, but not when in recording mode, the dither powers are likely not on. Are the camera settings correct for the seeing conditions and flux levels?

Back to Main Menu


Restarting Servers


Restarting Servers using the bootlaunch paradigm


A number of servers use an interim bootlaunch paradigm to restart. This is confined to servers that run on ubuntu machines, namely the telescope bunker computers and gps. The basic syntax is "bootlaunch_<server>" where "<server>" is replaced by the server the script is designed to address. The scripts have a number of safeties built in, so it is safe to run them even if a server is already running -- they just output the process ID of the running server. The scripts also take care of the entry in socket manager as well any serial port lock files. All the pertinent information is world writeable, so one should be able to run a bootlaunch script as observe.

One thing of note about the output of the bootlaunch scripts, they call a number of other programs which themselves have output that may be misleading in the context of bootlaunch. Chief among these is the output of "tsockman". If a server stopped unexpectedly, it may leave behind an entry in the socket manager. In order to launch a new server, one needs to clean out the socket manager entry if it is there. To do that, "tsockman remove <entry>" is called to remove "<entry>" before the new server is launched. If there is no entry, tsockman will respond with "Process by that name does not exist". THIS IS NORMAL and is not indicative of an error. The server in question launched (without fanfare) right after that output text.

Here are the available bootlaunch scripts as of June 2017:

gps computer:

  • bootlaunch_beamsamp -- Starts the beam sampler servers, BS1 and BS2.
  • bootlaunch_zaber -- Starts the ZABER_2 server.

telescope bunker computers:

  • bootlaunch_hut -- Starts the E1_HUT, E2_HUT, S1_HUT, S2_HUT, W1_HUT, or W2_HUT server, depending on the machine it's launched from.
  • bootlaunch_rpc -- Starts the RPC_E1, RPC_E2, RPC_S1, RPC_S2, RPC_W1, or RPC_W2 server, depending on the machine it's launched from.
  • bootlaunch_weather -- Starts the E1_WEATHER, E2_WEATHER, S1_WEATHER, S2_WEATHER, W1_WEATHER, or W2_WEATHER server, depending on the machine it's launched from.
  • bootlaunch_lower -- Starts the E1_Lower, E2_Lower, S1_Lower, S2_Lower, W1_Lower, or W2_Lower cylinder server, depending on the machine it's launched from.
  • bootlaunch_upper -- Starts the E1_Upper, E2_Upper, S1_Upper, S2_Upper, W1_Upper, or W2_Upper server, depending on the machine it's launched from.

Note: The bootlaunch scripts will not start a new server if there is an existing process running. Therefore, type "ps aux | grep server_name" where server_name is the name of the server. If there is a dead process running, look up the process identification number (PID) and type "kill -9 PID" to kill the process and then run the relevant bootlaunch script.

Restarting Servers using the rc.local file


This procedure is applicable to servers that have not switched over to the bootlaunch paradigm.

If a server is not running or Socket Manager reports that a server is dead, then look at the socket manager list to find out what computer the server runs on (socket_manager.list). You can also look at the up-to-date file by opening a terminal window and typing "less /chara/observe/socket_manager.list" Note that servers can be running fine, but if the Socket Manager drops the connection to them, they are as good as dead when it comes to functioning with other servers or as part of a larger sequence.

Log on to the relevant computer by typing the computer name (ctrscrut, ople, s1, ...). If the shortcut doesn't work then type "ssh name" where name is the computer name.

Find out if the server is running by typing "ps aux | grep server_name" where server_name is the name of the server.
[ctrscrut:599] ps aux | grep pico_1
observe 9281 0.0 0.0 61156 692 pts/3 S+ 13:58 0:00 grep pico_1
observe 12578 0.0 0.0 24524 11212 ? S Jun16 33:14 /usr/local/bin/pico_server /dev/ttyC8 /ctrscrut/chara/etc/pico_1.cfg

If the entry for the dead server shows up in the process list, then identify the process identification number (12578 for the example above) and kill the server by typing "kill -9 PID" where PID is the process identification number.

Look up the commands to restart the server by typing "more /etc/rc.local" (this is relevant for servers that run in the background). Press the space bar to scroll through the contents of the rc.local file. Locate the commands relevant for the server that needs to be restarted and copy and paste into a terminal window:

#Start PICO server for PICO #1
/bin/rm -f /var/lock/LCK..ttyC8
/usr/local/bin/tsockman remove PICO_1
/usr/local/bin/pico_server /dev/ttyC8 /ctrscrut/chara/etc/pico_1.cfg &

The first command removes the lock to allow the server to restart. The second command removes the name from the socket manager listing. The last command restarts the server. Note that if you are restarting the servers as observe, you will need to remove the part of the command in the rc.local file that saves information in /var/log/server_name.log file (the actual command typed should resemble the last line above).

There are text files on the desktop with many of the restart commands. Use these files for quick access to the relevant commands. The commands are edited and can be copied exactly as written. Files include Dome servers and all servers running on ctrscrut. Many of these commands are also located on the Restarting Servers page.

Shutters Server


The Shutters server can become unresponsive or disconnected from the Socket Manager. This server must be restarted from the lab and not from the Control Room. Follow these instructions to restart it. Note that Shutters runs on ople, not ctrscrut.

To start the shutter server on ople:

Log into the ople computer and kill the process labeled shutters with the PID as described in Restarting Servers above.
Turn off the power to the Shutters with the switch on the computer rack which is to the left of the computer desk and marked "SHUTTERS". Restart the Shutters server with the commands below. After restarting the server and testing the gui to see that it works, turn the SHUTTERS power back on with the switch. There is a printed sheet of directions in the lab to help you.

/usr/local/bin/tsockman rm shutters
ctrscrut/usr/local/bin/shutter_server /ctrscrut/chara/etc/shutter.cfg &


Back to Main Menu


Telescopes and Dome Servers


Here we discuss things that can go wrong with the telescopes.

The Telescope won't move or stopped moving


Have the powers been turned on to the drives? Are the scopes disabled? The usual state of the telescopes is disabled until enabled. This is due to the stall function of the scopes which eventually disables the scopes when they are stowed. Enable the scopes by hitting [ENABLE] on the dome gui or in the telescope gui control tab. If a scope disables itself during a slew, it may be just an overcautious stall function. Hit ENABLE and the scope should continue to slew. If it disables again, there may be something wrong at the scope, ie. a hatch left open, a ladder not put away, a tool on the floor, or something else physically impeding the motion of the scope. You will need to go to the dome to see what it is. The computer in the dome will give you control of the scope to turn it away from the problem.
Sometimes the dome guis get hung up and can cause erratic motion or no motion of the scopes. Check them for current times and continuous updates of numbers. If they are not updating, close the gui and open a new one. If a new one does not open, the dome server may be dead or Sockman lost track of it. See Dome Server Restart below.

The Telescope won't track


If a bright star is manually found and not slewed to by number, (usually after the scope has gotten lost), the telescope will not track as it does not know it has gone to a star. When you init the EL and AZ position of the scope to the star's position, you must enable the tracking by going to the dome gui, selecting the RA/DEC tab and hitting the [GO RA/DEC] button. The scope will track, but the pointing model may still be off. Try slewing to the same star and initing on it after it is locked in Tiptilt. This can be done in the dome with the computer at the telescope.

The Telescope won't stop moving or moves beyond the commanded position


Disable the scope whenever its motion is outside of what it is supposed to do. Turn the powers off to the axes as well on the Power gui. The dome server is usually to blame and will need to be restarted.

Dome Server Restart


To manually start the dome server:

1. Make sure the power to the drives is OFF.
2. Login to the relevant computer as root. For example, type "s1" or "ssh s1" to log on to S1.
3. Work out the process ID number (PID), either with the command

(s1:1001) tsockman get dome_S1
Name : dome_S1
Machine : s1.chara-array.org
PID : 29953
Commands : -1
Data : -1
Message : 4002
Restart : /usr/local/bin/dome_server -A33.7441 S1

or with
(s1:1003) ps aux | grep dome
theo 4473 0.0 0.0 61188 748 pts/3 S+ 10:45 0:00 grep dome
observe 29953 18.5 0.4 35596 9860 ? Sl Apr21 416:11 /usr/local/bin/dome_server -A33.7441 S

It can also be found by pulling up the LIST on SOCKMAN and selecting the relevant dome.

So in this case the PID is 29953.

4. Try and stop the server gracefully: kill -2 29953
5. You should then check that the server has indeed stopped:

[s1:600] tsockman get dome_S1
Name : dome_S1
Machine : s1.chara-array.org
PID : 15635
Commands : -1
Data : -1
Message : 2008
Restart : /usr/local/bin/dome_server -A41.0166 S1

If the socket manager still thinks it's running you will need to stop it forcefully: kill -9 29953; tsockman rm dome_S1

6. Restart the dome server by copying the command at the end of the /etc/rc.local file:

[s1:602] more /etc/rc.local
<<< Press space bar or enter to scroll through file>>>

#Run the dome server

/usr/local/bin/tsockman remove dome_S1
/usr/local/bin/dome_server -A41.0166 S1 &

(Note: the part of the command that saves information to /var/log/dome_S1.log has been removed.)

7. Turn the power to the drives back on.
8. Hit REOPEN and ENABLE on the domegtk, and type "otcs" in the telescope server.

You may have to reinitialize the scope on a bright star. If the powers were turned off quickly when the problem was noticed, the position of the scope should be retained and slewing to a bright star will get it in the finder. If not, you may need to go out to the telescope to find the bright star to reacquire the scopes position.


Telescope clock is not correct


The NTP server has obsoleted this problem and fix as of 2014.

Telescope is not receiving the commanded position for a target


Sometimes it happens that a telescope receives the wrong position for a target or does not receive the commanded position at all. The commanded position is listed on the telescope server in the first column under TCS Az/El; the second column lists the actual position of the telescope. Try entering the star designation directly into the telescope server, ie. hd 123456. If it does not accept the number, try closing and restarting the telescope server and hitting repoen on Cosmic Debris and the telescope gui. Try entering the star into the server again. If that does not work, it is possible that something is wrong with the dome server. To restart the dome server follow these steps:
  • DISABLE the telescope using either the telescope or dome GUI.
  • Turn off the power for the telescope (both AZ/EL).
  • First shutdown the telescope server
  • Then use SOCKMAN to select the appropriate dome server from the list (dome_E1, etc) Get the PID number for reference.
  • The dome server will need to be restarted from the command line in a terminal.
  • Open a terminal window and log on to the telescope computer by typing "s1", "s2", "e1", "e2", "w1", or "w2"
  • Look for instruction in the /etc/rc.local file by typing "more /etc/rc.local". Scroll through the file by hitting space bar or Enter.
  • Locate instruction listed under "#Run the dome server". Copy and paste the commands at the prompt directly on the telescope computer:
    • /usr/local/bin/tsockman rm dome_S1
    • /usr/local/bin/dome_server -A33.7441 S1 >& /var/log/dome_S1.log
  • (Replace S1 with commands for appropriate telescope)
  • After the dome server is running, re-open the telescope server and click [REOPEN} on the dome and telescope GUIs and on Cosmic Debris.
  • ENABLE the telescope using the telescope or dome GUI. Check to see if the command and telescope positions are correct. If they are, then turn on the power for the telescope drives. Re-enter the star information by entering the HD number in the telescope server and click [GO NEXT] to send the telescope to the star. Make sure that the telescope is behaving as expected. The telescope might have drifted a bit, so if you can't find the star, it might be necessary to use the Telrad on a bright star to re-initialize the pointing.
  • Restart commands for the dome servers are also listed on the desktop in a text file.


Telescope is tracking poorly, overshooting in slew, oscillating.


This might mean that the gain for the tracking servo is wrong. Note that changing this gain can be dangerous, especially if you set it too high as that can cause the telescope to oscillated and damage the drives. Please only do this if you are very very sure that it is necessary. Symptoms of bad gain are:

The scope over shoots the position while slewing. The star will be seen to move out of the window and may come back after a few seconds. This means the slewing gain is too low.
The scope oscillates when tracking or after a slew. The star will be tracing an ellipse, figure eight or other looping shape. This means the tracking gain is too low. You can damp this out with the telescope or dome gui by disabling the scope, then re-enabling it. Adjust the gain upward and watch it on the next slew.

In all cases if either gain is too high the scope will go into "Fog Horn" mode, which is bad. You always want to use the lowest gain that still allows the scope to work as best as possible. If the tiptilt tells you the scope is oscillating slowly, the gain may be too low. If it is oscillating quickly it may be too high.

On 10-22-2016, the gain settings were:

Scope
AZ Slewing
EL Slewing
AZ Tracking
EL Tracking





S1
4
4
13
13
S2
7
7
10
10
E1
7
7
13
10
E2
7
7
13
10
W1
13
10
13
10
W2
7
7
13
13

Note that these values may change as the temperatures change rapidly. These are still being adjusted so this is by no means final. Be sure to set mode back to AUTO if changing the gains left it in Slewing or Tracking mode.

Back to Main Menu


Dome issues


At times the domes do not rotate, open or close, or otherwise behave. Some problems are simple and others are more complex.

Dome does not rotate


Sometimes when observing, the dome will not follow the telescope during a slew. This can happen when the Autodome feature is not turned on. Click the ON button on the MAIN tab of the telescope gui to enable it. This may happen after a server restart so always check the dome position with the spycam during a slew after a server restart. Also make sure the target position of the dome matches the telescope's position. If not, it will insist on being in the wrong place. If it is not at the same AZ as the scope, manually move it until it is centered on the telescope in spycam 1. If the dome AZ does not read the same as the telescope AZ, enter the scope AZ in the position box of the DOME tab of appropriate dome server and hit the INIT POS button to tell it at what AZ it is.

If the dome does not turn at all, even with the manual controls on the telescope or dome guis, the control may be set to manual on the control box. This can happen if there was work done at the dome during the day. If the dome opens, but does not turn, check to see that control of the dome rotation is in the computer position and not manual on the dome rotation controller box just inside the door of the bunker.

Dome does not open


If the dome won't open, try hitting the SLIT CLOSE button on the telescope gui first. It may be that there was a computer or server issue and the computer thinks the dome is open already and won't allow it to open again. Hit the SLIT OPEN button to see if it works now. If it still won't open, go to the bunker and look to see if the power switch is on to the dome. It will be up and red if on.
Also look to see if the power cable is connected to the computer box. If it's connected up high to the manual dome controls, the cable needs to be brought back to the computer box connector. Turn the power off first to the dome before disconnecting the power cable. Slide it onto the connector at the box and turn the collar to lock it in place. Turn the power back on and see if you can open the dome from the computer in the bunker.

Dome does not close


If a dome will not close after observing, it may need to be closed manually at the bunker. Turn the power to the dome off at the switch, disconnect the cable from the computer control box and connect it to the manual control switches above. Turn on the power to the dome after connecting the cable and use the two controls to close the dropout first, then the slit. The dropout must be fully closed and the slit closed over it to be shut properly. Inform the staff that the dome did not function properly so someone can look at it in the morning.


HUT servers


I can't change the camera settings on the TV


The HUT servers control functions such as finder and acq exposure times and gains, heater and dehumidifier usage, and various AO functions. An observer may find that the camera settings do not display or are not adjustable. The HUT server may be the cause if it has quit. To see if it is the server, open the HUT gui for the desired telescope from the CHARA menu. If the camera settings are displayed and the settings can be changed from the gui, then the HUT server is ok. Restart the telescope server to reopen the connection to the HUT server and hit reopen on Cosmic Debris as well. If the gui is not functioning, the HUT server will need to be restarted. The HUT servers run on the computer at each telescope. Find the PID from Sockman for the server you want to shut down. Open a terminal and log into the correct telescope computer (for example, type "s1" to log on to S1). Use the command "ps aux | grep HUT" to find the process you need to shut down. The number should match the PID from Sockman. Kill the process by typing "kill -9 pid" where pid is the process number and remove the entry from Sockman with the REMOVE button on the gui or by command line. Restart the hut server using the new "bootlaunch_hut" command described in the section on Restarting Servers. If the server won't restart, a reboot of the power supply in the telescope bunker might be necessary. The power supply that controls the acquisition and finder cameras as well as their controllers is located on top of the computer rack in each bunker. The power supply has green readouts of volts and current. After turning the power off for 10 seconds and back on, try restarting the server from the computer in the bunker to see if it starts cleanly. If so, then restart the telescope server, reopen the connection to the telescope gui, and hit REOPEN on Cosmic Debris. Part of the HUT server also controls the AO table. If the AOB part of the HUT server doesn't work, then the power supply on the back of the AO table in the telescope dome might need recycling. This power supply controls the actuators at M2 and the AO table. The power supply box is a 6x9 inch aluminum box on the back of the AO table, behind the keyboard and monitor. Turn it off with the power button on the bottom edge of the box, wait 5 seconds and turn it back on. The HUT server should now restart cleanly. Restart the telescope server as well to make the connections to the telescope gui. Hit REOPEN in Cosmic Debris if you are observing to make all needed connections.


Tiptilt Server


The tiptilt server controls the CCD based tiptilt detection system.

Before you start the tiptilt server, you must ensure that the power to the cooling system and the CCD iteslf is on. It is extremely important that the cooler be running before you turn on the CCD and is only turned off if you are sure the CCD is NOT running. You can start the server from the X windows menu or with the command xtiptilt.

Note that there are background counts and read noise to deal with. Whenever you change the frame rate, please ensure that the bias frame is OK. The server will attempt to load an old bias frame that should work, but if things are not working, try making a new bias frame by ensuring that the detector is in the dark and typing "mkbias" into the tiptilt server.

In the tiptilt GUI windows, the white dots represent the starlight while the green dots represent the motion applied to telescope's secondary mirror to keep the starlight centered. When tiptilt is locked the white dots will be brought to the center of the tiptilt window. The green dots should be mostly centered also. W2 and E2 telescopes have a small oscillation that show as back and forth plots of the green dots.


Tiptilt server complains about the CCD


Is the CCD turned on? When the tiptilt server starts up it tries no more than five times to communicate with the CCD. If they all fail, it will give up. If this happens, try cycling the power to the CCD and try again. If this fails, connect to the tiptilt machine and type the command rtccdAPIDemo, which should return with no errors. Try this command a few times, but if it still fails, there is a more serious problem. Turn off the CCD and reboot the tiptilt computer. If it still fails, I am afraid you are in more serious trouble.

Note that it is never a good idea to reboot machines unless you are very very sure it is necessary. The only reason to reboot tiptilt, other than a lock up of some kind, is that the clock interrupt has failed. You can test this by running the command "testclock" on a tiptilt command line. If this says the clock is working do not reboot the machine.

Also note that cycling the power on the CCD can cause harm so be sure you need to do it before trying it. Also, it is important to wait for at least 20 seconds after turning off the power before turning it on again.


Tiptilt doesn't seem to be talking to the telescopes


Sometimes the telescope server will not show that TT is running. It will show 0Hz for a signal rate for TT. Running TIPTILT COMM will not get it started while other scopes do show it starting. Close and restart any telescope servers that won't connect after two tries of TIPTILT COMM.
Note: There is more info in the software manual on this topic, but I wasn't sure if it was still relevant.


Tiptilt server says the clock isn't running

This is a bug, that I hope to fix soon, in the real time part of the CCD code. It is caused by the clock in the tiptilt system either not running at all or having been set to a time very different from the last time the CCD ran. For the time being the only solution is to reboot tiptilt, but do so from the lab and go into the BIOS. Make sure that interrupt 11 has been set to ISA legacy, save the BIOS and reboot. Also, sometimes Serial Port 3 grabs IRQ 11 which stops the clock from running. Since there is no serial port 2 it's safe to disable this in the BIOS. This problem normally comes up when there has been a power outage.

Sometimes syncing the clock can also cause this problem, but that should be fixed soon. If it does, exit the tiptilt server, log in as root, and reload the tiptilt model using the following commands:
/sbin/rmmod tiptilt_rt
/sgin/insmod /usr/local/modules/tiptilt_rt.o

Note that it is never a good idea to reboot machines unless you are very very sure it is necessary. The only reason to reboot tiptilt, other than a lock up of some kind, is that the clock interrupt has failed. You can test this by running the command "testclock" on a tiptilt command line. If this says the clock is working do not reboot the machine.

Tiptilt is not locking on a star or locks, but lets the star drift away

  • Are all the mirror covers open?
  • Has TIPTILT COMM been run from Cosmic Debris?
  • Check the ACQ alignment to make sure the tick marks are centered on the laser.
  • If there is plenty of starlight getting into tiptilt, then try re-initializing by clicking [TIPTILT COMM] on Cosmic Debris.
  • If there are a large number of background counts on TT, then close the M5 cover on the telescope, and click [DBIAS] on the TIPLTILT GUI to clear the background counts. Re-open the M5 cover.
  • If a star drifts even with TT locked, there could be a bright sky or light from other beams getting into the affected telescope's beam. This will show in the TT server as much lower counts than the other locked stars, but still high enough to lock TT. Use the laser to find the correct position to lock TT.
  • If TT unlocks and the star drifts, the TT servo may not have engaged. This is engaged with the TIPTILT button on the telescope gui. Clicking it may not start the servo the first time. Try it again if the green dots drift on the TT plot windows or if the Servo status in the telescope gui reads None instead of Wobb 1.
  • Are the TIPTILT buttons turned on from the POWER GUI?

Tiptilt servo oscillates

You will see the oscillation in the green dots of the tiptilt GUI windows. Sometimes you can also see the oscillation in the white starlight dots or as an elongation of the star when looking at the ACQ field while tiptilt is locked. Some scopes have an oscillation that has not yet been diagnosed. W2 is one that usually oscillates. A diagonal motion in the tiptilt box indicates an oscillation in one axis only, while a vertical/horizontal motion indicates an oscillation in both directions. Motion from the upper right to lower left corresponds to elevation axis while motion from the upper left to lower right corresponds to the azimuth axis. (I think you can check direction by typing sin into telescope server to send sine waves to the telescope.) There are a few ways you can try to correct the oscillation manually tuning the servo:
  • Type "tune" into the tiptilt server. Select the appropriate telescope. Turning the gain down normally helps.
  • Type "tune" into the appropriate telescope server. The default value for the proportional term is -0.5 and differential term is 0.0. Adjust these values between -0.2 to -1.0 for proportional and 0.0 to 0.2 for differential.
  • Read section on adjusting the Telescope Tracking Gain using the Dome Server GUI (ref).

Tiptilt is saturating

  • Tiptilt saturates at ~ 200,000 counts. If you are near this limit, you can reduce the TT exposure time to lower the number of counts.
  • Set the TT exposure time on Cosmic Debris in the box for "Tiptilt (mS)". This will set the TT exposure time when slewing to a new target.
  • To change the exposure when already at a target, then click the [EXP] button on the TipTilt GUI. This will bring up a dialog box where you can enter the new exposure time in msec. Check to make sure the tiptilt frequency (in Hz) changes on the tiptilt server after changing the exposure time. If the frame rate doesn't change, then set the exposure time back to the old value and try entering the new value again. You might have to do this a few times to actually get the frame rate to change.
  • Frame rates for given exposure times:
    • ExpTime = 5 msec, Frame Rate = 157 Hz
    • ExpTime = 2 msec, Frame Rate = 299 Hz
    • ExpTime = 1 msec, Frame Rate = 427 Hz

Tiptilt counts are way too low

Try using a slower frame rate or increasing the NSUM. Also ensure that the acquistion is properly aligned with the laser. To change the frame rate, click the [EXP] button on the tiptilt GUI and enter a longer integration time. Remember to change "Tiptilt (mS)" on Cosmic Debris to keep the same exposure time when slewing to the next target.

Tiptilt counts are negative

The bias frame is bad. Get a new one or turn it off.

Back to Main Menu

OPLE and Metrology


OPLE Server doesn't open - complains about not finding run_ople

This usually happens if someone runs a "make clean" in the cvs tree but doesn't follow up with a "make install" on ople. To generate the run_ople file, follow these steps:
  • Open a terminal window and type "ople" to log onto the ople machine.
  • Change to the appropriate directory under the CVS tree by typing "cd control/cliserv/ople"
  • Update the CVS tree by typing "cvs update"
  • Run the Makefile script by typing "make"
  • Log on as root by typing "su" and entering password.
  • Type "make install"
  • This will install run_ople. Try opening the ople server again.

Metrology and Homing carts

  • Good metrology signals are important to the proper positioning of the carts. Monitor the signal strength by running [RUN MULTIPLE] on the Metrology Monitor. Place the windows above the TV windows for each scope you are using. They should show white sine waves that are around the height of the window. Erratic, fluctuating waves indicate self-interference or a weak signal. This may cause the carts to lose their place as the signal strength falls too low. Red waves indicate that some displayed signal has gone too low and the carts will all need to be homed. A careful adjustment of the MET2 mirror can sometimes bring the signal back. Do not adjust the MET1 mirror.
  • To home the carts, turn off the [OL] and [MAN] buttons on each cart and it will automatically return to the front switch. If the cart has no issue, it will arrive at the target position of 0m and the home switch at the same time. The X in the OPLE server under the HM will indicate it has homed to the home switch. If a cart does not reach the home switch when it returns to position 0m, it was lost and likely the cause of the difficulty in finding fringes. Hit the [HOME] button and it should move forward and trigger the home switch. Hit [TRACK] to home it to the new home position.

Back to Main Menu

Beam Samplers


Beam Samplers are not moving

This happens most frequently with the E1 beam sampler. If you are doing alignments and are using E1, try to move it forward first to see if it dies. If it does, follow this procedure.
  • Turn off power on Newport Universal Motion Controller/Driver for the Beam Samplers (on top of rack in computer area, labelled "BS")
  • Then go behind the racks and locate the serial port for BS. Look for all of the blue ethernet cables coming out. BS is labelled in prt 2. Remove the ethernet cable for BS. The step stool is available to make it easier to reach the blue ethernet cable at the back of the BS controller box.
  • Restart BS controller (press power button again). Wait for all axes to appear on screen.
  • Plug the ethernet cable back in.
  • Restart BS server on ctrscrut using commands in /etc/rc.local (see instructions below)
  • Reopen Beam Sampler GUI. You will need to home each Beam Sampler one by one. You should be able to do this from the regular Beam Sampler GUI. If that doesn't work, try opening "espgtk BS" from the command line.

Restarting the Beam Sampler Server

The most common cause for getting multiple copies of a server is using the socket manage RESTART button, which is flaky at best. Don't do this. If you need to restart a server you should do it manually.

1. Make sure you are logged into the right machine:

ssh ctrscrut

2. If you are not sure see if the socket manager will tell you. If it doesn't have a look in the file:

/ctrscrut/chara/etc/sock_manager/socket_manager.list

3. See if there are any ghosts running

ps aux | grep esp_server

Yes, it's hard to know which one it is with the esp servers. You can work out which ones are ghosts by typing the command

tsockman |grep ctrscrut

which will give you a list of the servers running on the machine you are interested in. Checking for non-matching PIDs will tell you which processes you need to stop.

4. Stop those processes:

kill -9 PID1 PID2 ....

5. Make sure there is no sign of it in the socket manager

tsockman rm BS

5. Restart the beam sampler servers (this starts both beam sampler servers BS1 and BS2):

bootlaunch_beamsamp


"Failed to request position of S1"


Cosmic Debris reports "Failed to request position of S1" when trying to set the beam order. This indicates that ople is no longer talking to the beam samplers. If the beam sampler server has been restarted recently, then the ople server will also need to be restarted to re-establish the connection.

The beam sample server runs on ctrscrut. The ople server tries to open a connection to the beam sampler when it starts. If the beam sampler is not there, or dies,
you need to restart the ople server as there is no command to reopen that connection.

Back to Main Menu

Remote Observing


VPN connection is not working

To re-establish the VPN connection, follow these steps:
  • Open VPN status page from the firefox web broswer. To do this, click on the "AROC VPN reconnect" icon on the zoot desktop. Alternatively, you can open the firefox web browser and go to 192.168.3.1/vpn_sta.htm
  • The browser will prompt you for the admin user name and password. If these entries are not already filled in, then look for this information on the papers near the console.
  • If the browser does not open, then check if there are other instances of firefox already running. To do this, open a terminal window, type "ps auxw | grep firefox", and kill the other instances of firefox so that you can open a new window.
  • When the VPN connection is working, it will report "Phase 1: M-ESTABLISHED / Phase 2: ESTABLISHED" under the state column.
  • If either Phase 1 or Phase 2 states say IDLE for a given VPN, then you need to re-establish the connection. To do this, click on the "DROP" button for that VPN. Then click on the "CONNECT" button and wait for the Phase 1 and Phase 2 connections to be established.
  • Make sure to close Firefox by using CTRL-Q to quit the program and not just close the window.

Back to Main Menu

General Problems


GUIs are frozen

Check to see if Sockman is working. Click the [LIST] button on sockman. If the list doesn't come up, Sockman may be hung up. Follow the Sockman restart procedure in the text file on the desktop if it won't respond.
Hit [REOPEN] on the gui to see if it reconnects with its server. if that does not fix it, the server may be dead. From the LIST on Sockman, select the server suspected of being frozen. Ping it and if it reports back as being dead, see Server is frozen section below.
Telescope guis often freeze. Hit reopen on the gui to bring it back to life. Hitting [MOVE] on the telescope gui before the star has stopped moving can cause the gui to freeze. Avoid doing this. Close the dead one and open a new one to fix the problem.

Server display is blank or filled with jibberish

  • If the top of a server goes blank, try typing "sb" (start background) on the server command line.
  • If the server screen fills with jibberish, try hitting CTRL-l in the server to clear it.

Server is frozen

Try quitting the server using CTRL-C and entering "Y". If that doesn't work, try the following:
  • On Sockman, click [LIST] and select the appropriate server. It will give a PID number you can use to kill the server in a terminal. Open a terminal and follow the directions for Restarting Servers above.
  • If a server window does not close, click the X in the upper right corner to close it, but only after getting the PID from Sockman.
  • Click [REOPEN] on Cosmic Debris and relevant GUIs to re-initialize communication with the server after it is restarted.
  • A folder is on the desktop that has the restart commands for CTRSCRUT servers and the Shutters server restart command running on OPLE. Use it to restart servers that will not reopen from the menu.

PAVO Server - Error communicating with IFW

The following error message sometimes appears when starting the PAVO server:
"Error communicating with IFW. Could not read from IFW. Is the filter wheel plugged in and on?"
Most of the time, restarting the PAVO server will clear this error message and allow PAVO to communicate with the filter wheel. If restarting the server doesn't help, then check that the cables going into the small, black IFW box on the PAVO table and into the back of the PAVO computer are plugged in securely.

Adding or Finding a Star in the CHARA Database

DBADD is a command entered in a terminal window that allows you to add a star to the CHARA database or look up a name of a star, nova or AGN to see if it already has a CHARA number. Valid requests can look like dbadd Vega or dbadd HD 123456. Identifiers for a target can be a common name, an IRC, HR, HD, SAO, FK5, HIP, GJ, or 2MASS designation. If a star, AGN or nova is in the database, it will return a CHARA number.

Targets that are not in the CHARA database

If a star, nova or AGN is not in the CHARA database, dbadd will query the SIMBAD database to see if it can be found there. If it returns a result, the target's coordinates, magnitudes, and spectral type will be given. If this matches your object, then accept the result and a new CHARA number will be issued for the object.
You can also try to find another identifier for the star by using the
SIMBAD database: http://simbad.u-strasbg.fr/simbad/sim-fbasic

SIMBAD will also show many other identifiers for your star and most likely will show one of the 8 database identifiers that Cosmic Debris recognizes. If you cannot find an identifier that Cosmic Debris or SIMBAD recognizes, you must enter the object manually with the instructions below. Examples of objects not identified are AGN's or novae.

Targets that are not in the CHARA database or in SIMBAD

Some objects like new novae or AGN's are not well studied and therefore are not in SIMBAD. Entering dbadd <star designation> for a star not in the SIMBAD database returns the message "Simbad is unable to find an object matching <star designation> Try using a different catalog designation, or use the "-m" switch to add the object manually."

The command dbadd -m is used by entering the star name and coordinates in this format:

Usage: dbadd -m <name> <RA> <Dec>

<name>: Object ID (no spaces)

<RA> : XXhXXmXX.X or XXhXX.X (no spaces)

<Dec> : XXdXXmXX.X or XXdXX.X (no spaces)

There are cases where an object is not in SIMBAD and doesn't return a result in dbadd, but is in fact in the CHARA database. Novae and AGN's are the likely objects that cause this result. At times, objects have multiple entries and several CHARA numbers since the names can be so unique. A system will be developed to find these entries without knowing or having a conventional database designation.

Binary stars

Some binary stars have a common HD number with an A or B after them. These can cause problems as Cosmic Debris does not accept non numerical entries when filling in star designation. These stars are likely in CHARA's database but need to be searched in dbadd or on SIMBAD to get the CHARA number or another designation.
When a recognized identifier is entered, the HD number with A or B will usually display. Confirm that the coordinates, magnitudes and spectral type match the star desired. If they do not match, you may have the wrong star or the database may have the wrong info and the baseline solution will be affected. Offsets based on incorrect coordinates or misidentifications can move the fringes many thousands of microns away from the calculated position. This can happen when observing a dim companion (B) using the brighter (A) companion's coordinates. Inform the observers that noting and using the CHARA number will save time next time the target is observed.

Editing the database

If you find a mistake in the database, please send an email to Nils to have it corrected. Identify what you believe to be the error and what is the correct information.

Adding an anchor to the wiki page

To add an anchor to the wikipage, type the following where you want the anchor inserted (without the spaces between the symbols):
[ [ #AnchorName ] ]

Highlight the text you want to be a link to the anchor, click the link symbol above and select anchor. Select a page(such as Trouble Shooting) where the anchor is located and give the same name for the anchor.

Back to Main Menu

Last updated 2017-04-12
grep bs /etc/rc.local

6. Run it

/usr/local/bin/esp_server /dev/ttyC1 bs &


"Failed to request position of S1"


Cosmic Debris reports "Failed to request position of S1" when trying to set the beam order. This indicates that ople is no longer talking to the beam samplers. If the beam sampler server has been restarted recently, then the ople server will also need to be restarted to re-establish the connection.

The beam sample server runs on ctrscrut. The ople server tries to open a connection to the beam sampler when it starts. If the beam sampler is not there, or dies,
you need to restart the ople server as there is no command to reopen that connection.

Back to Main Menu

Remote Observing


VPN connection is not working

To re-establish the VPN connection, follow these steps:
  • Open VPN status page from the firefox web broswer. To do this, click on the "AROC VPN reconnect" icon on the zoot desktop. Alternatively, you can open the firefox web browser and go to 192.168.3.1/vpn_sta.htm
  • The browser will prompt you for the admin user name and password. If these entries are not already filled in, then look for this information on the papers near the console.
  • If the browser does not open, then check if there are other instances of firefox already running. To do this, open a terminal window, type "ps auxw | grep firefox", and kill the other instances of firefox so that you can open a new window.
  • When the VPN connection is working, it will report "Phase 1: M-ESTABLISHED / Phase 2: ESTABLISHED" under the state column.
  • If either Phase 1 or Phase 2 states say IDLE for a given VPN, then you need to re-establish the connection. To do this, click on the "DROP" button for that VPN. Then click on the "CONNECT" button and wait for the Phase 1 and Phase 2 connections to be established.
  • Make sure to close Firefox by using CTRL-Q to quit the program and not just close the window.

Back to Main Menu

General Problems


GUIs are frozen

Check to see if Sockman is working. Click the [LIST] button on sockman. If the list doesn't come up, Sockman may be hung up. Follow the Sockman restart procedure in the text file on the desktop if it won't respond.
Hit [REOPEN] on the gui to see if it reconnects with its server. if that does not fix it, the server may be dead. From the LIST on Sockman, select the server suspected of being frozen. Ping it and if it reports back as being dead, see Server is frozen section below.
Telescope guis often freeze. Hit reopen on the gui to bring it back to life. Hitting [MOVE] on the telescope gui before the star has stopped moving can cause the gui to freeze. Avoid doing this. Close the dead one and open a new one to fix the problem.

Server display is blank or filled with jibberish

  • If the top of a server goes blank, try typing "sb" (start background) on the server command line.
  • If the server screen fills with jibberish, try hitting CTRL-l in the server to clear it.

Server is frozen

Try quitting the server using CTRL-C and entering "Y". If that doesn't work, try the following:
  • On Sockman, click [LIST] and select the appropriate server. It will give a PID number you can use to kill the server in a terminal. Open a terminal and follow the directions for Restarting Servers above.
  • If a server window does not close, click the X in the upper right corner to close it, but only after getting the PID from Sockman.
  • Click [REOPEN] on Cosmic Debris and relevant GUIs to re-initialize communication with the server after it is restarted.
  • A folder is on the desktop that has the restart commands for CTRSCRUT servers and the Shutters server restart command running on OPLE. Use it to restart servers that will not reopen from the menu.

PAVO Server - Error communicating with IFW

The following error message sometimes appears when starting the PAVO server:
"Error communicating with IFW. Could not read from IFW. Is the filter wheel plugged in and on?"
Most of the time, restarting the PAVO server will clear this error message and allow PAVO to communicate with the filter wheel. If restarting the server doesn't help, then check that the cables going into the small, black IFW box on the PAVO table and into the back of the PAVO computer are plugged in securely.

Adding or Finding a Star in the CHARA Database

DBADD is a command entered in a terminal window that allows you to add a star to the CHARA database or look up a name of a star, nova or AGN to see if it already has a CHARA number. Valid requests can look like dbadd Vega or dbadd HD 123456. Identifiers for a target can be a common name, an IRC, HR, HD, SAO, FK5, HIP, GJ, or 2MASS designation. If a star, AGN or nova is in the database, it will return a CHARA number.

Targets that are not in the CHARA database

If a star, nova or AGN is not in the CHARA database, dbadd will query the SIMBAD database to see if it can be found there. If it returns a result, the target's coordinates, magnitudes, and spectral type will be given. If this matches your object, then accept the result and a new CHARA number will be issued for the object.
You can also try to find another identifier for the star by using the
SIMBAD database: http://simbad.u-strasbg.fr/simbad/sim-fbasic

SIMBAD will also show many other identifiers for your star and most likely will show one of the 8 database identifiers that Cosmic Debris recognizes. If you cannot find an identifier that Cosmic Debris or SIMBAD recognizes, you must enter the object manually with the instructions below. Examples of objects not identified are AGN's or novae.

Targets that are not in the CHARA database or in SIMBAD

Some objects like new novae or AGN's are not well studied and therefore are not in SIMBAD. Entering dbadd <star designation> for a star not in the SIMBAD database returns the message "Simbad is unable to find an object matching <star designation> Try using a different catalog designation, or use the "-m" switch to add the object manually."

The command dbadd -m is used by entering the star name and coordinates in this format:

Usage: dbadd -m <name> <RA> <Dec>

<name>: Object ID (no spaces)

<RA> : XXhXXmXX.X or XXhXX.X (no spaces)

<Dec> : XXdXXmXX.X or XXdXX.X (no spaces)

There are cases where an object is not in SIMBAD and doesn't return a result in dbadd, but is in fact in the CHARA database. Novae and AGN's are the likely objects that cause this result. At times, objects have multiple entries and several CHARA numbers since the names can be so unique. A system will be developed to find these entries without knowing or having a conventional database designation.

Binary stars

Some binary stars have a common HD number with an A or B after them. These can cause problems as Cosmic Debris does not accept non numerical entries when filling in star designation. These stars are likely in CHARA's database but need to be searched in dbadd or on SIMBAD to get the CHARA number or another designation.
When a recognized identifier is entered, the HD number with A or B will usually display. Confirm that the coordinates, magnitudes and spectral type match the star desired. If they do not match, you may have the wrong star or the database may have the wrong info and the baseline solution will be affected. Offsets based on incorrect coordinates or misidentifications can move the fringes many thousands of microns away from the calculated position. This can happen when observing a dim companion (B) using the brighter (A) companion's coordinates. Inform the observers that noting and using the CHARA number will save time next time the target is observed.

Editing the database

If you find a mistake in the database, please send an email to Nils to have it corrected. Identify what you believe to be the error and what is the correct information.

Adding an anchor to the wiki page

To add an anchor to the wikipage, type the following where you want the anchor inserted (without the spaces between the symbols):
[ [ #AnchorName ] ]

Highlight the text you want to be a link to the anchor, click the link symbol above and select anchor. Select a page(such as Trouble Shooting) where the anchor is located and give the same name for the anchor.

Back to Main Menu

Last updated 2017-04-12